OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

Ok, so we know that the slope has to be equal to 2. So check. Now, lets set up a simple equation with the formula for slope. So... y - 5/ x- 3 = 2 y - 5 = 2(x-3) y - 5 = 2x - 6 y = 2x -1

OpenStudy (anonymous):

Btw, if you're not getting what he said, he use point-slope form \[y - y_{1} = m(x - x_{1})\]